Company X wants to increase its capital to finance the international expansion included in its strategic plan for the period X6-X9. To do so, the company is considering the incorporation of a new shareholder that subscribes 80,000 new shares.

Equity for Company X before this capital increase is:

Capital 3,000,000 (300,000 shares, €10 par value each)

Reserves 1,200,000

Current shareholders do not wish that this capital increase damages the net book value of their current shares.

Required

Can you record the corresponding entry/ies for this capital increase knowing that the new shareholder will pay out the legal minimum required at the same moment as the capital increase, and that he/she will pay the rest within the next two months?

Questions:

1. What is the net book value of a share of Company X before and after the capital increase?

2. Discuss the difference between book value and market value



1
Expert's answer
2021-03-04T10:41:03-0500

Solution:

Journal entries:

We assume the company has sold 80,000 shares at €15 per share. As per the terms of the issue €12 has been received on March X6, while the remaining amount will be received in full on May X6.

Journal entry for the transaction will be as follows:

                                                                                                       Debit                      Credit

March X6        Bank (80,000 shares X €12 per share)           960,000                     

                          Share Capital (80,000 shares X €10 per share)            800,000                               

                         Paid-In Capital in Excess of Par Value

                         (80,000 shares X €2 per share)                                                                    160,000

(To record the issuance of 80,000 shares of stock banked at €12)

 

 

May X6             Bank (80,000 shares X €3 per share)                240,000

 

                          Paid-In Capital in Excess of Par Value

                          (80,000 shares X €3 per share)                                                                   240,000

                          

(To record the balance of 80,000 shares issued at €3 received)

 

1.). The net book value of a share of company X before the capital increase.

NBV of a share = Shareholders Equity/Shares Outstanding

Shareholders Equity = Capital + Reserves = 3,000,000 + 1,200,000 = 4,200,000

Shares Outstanding = 300,000 shares

                               = 4,200,000/300,000

                               = €14

The netbook value of a share of company X after the capital increase.

NBV of a share = Shareholders Equity/Shares Outstanding

New capital = 3,000,000 + 1,200,000 = 4,200,000

New shares outstanding = 300,000 + 80,000 = 380,000 shares

Shareholders Equity = Capital + Reserves = 4,200,000 + 1,200,000 = 5,400,000

Shares Outstanding = 380,000 shares

                               = 5,400,000/380,000

                               = €14.21

 

2.). The difference between book value and market value:

The book value is the net value of a company’s assets found on its balance sheet or the shareholder's equity and it is basically the total amount all shareholders will get if the company is liquidated. The book value equals the difference between a company’s total assets and total liabilities.

A market value, on the other hand, is the value of a company based on the financial markets. It is calculated by multiplying the current stock price by the number of outstanding shares that are trading in the market. It is also known as market capitalization.